Hello,

I need help with these two problems. I dont know how to set them up, If could set them up for me then I would be happy to solve them.
The Divine Chocolate store want to create a new truffle mix that consists of milk chocolate and hazelnut. If hazelnuts sell for $2 per pound and milk chocolate sells for $4 per pound, how many pounds of each should they mix together to create a 18-pound box that sells for $3 per pound?
A chemist wants to mix a 60% alcohol solution with 9 liters of a 66% alcohol solution to produce a solution that is 62% alcohol. How many liters of the 60% alcohol solution should be used?
Thank you so much!

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

The Divine Chocolate store want to create a new truffle mix that consists of milk chocolate and hazelnut. If hazelnuts sell for $2 per pound and milk chocolate sells for $4 per pound, how many pounds of each should they mix together to create a 18-pound box that sells for $3 per pound?

let x = number of pounds of hazelnut.
let y = number of pounds of mixed chocolate.

you know that the total pounds has to be 18, so x + y = 18

you know that the total money to be earned is 3 dollars a pound times 18 pounds = 54 dollars.

you know that hazelnuts cost 2 dollars a pound and mixed chocolates cost 4 dollars a pound.

therefore 2x + 4y = 54

the two equations that you have to solve simultaneously are:

2x + 4y = 54
x + y = 18

work it out and see what you get.

don't forget to confirm if you answer is correct.
----------------------------------------------------------------------------

A chemist wants to mix a 60% alcohol solution with 9 liters of a 66% alcohol solution to produce a solution that is 62% alcohol. How many liters of the 60% alcohol solution should be used?

let x = number of liters of 60% solution.

x + 9 is number of liters of total solution.

your equation is:

.6 * x + .66 * 9 = .62 * (x + 9)

work it out and see what you get.

don't forget to confirm if you answer is correct.

------------------------------------------------------

let me know if you have any problems getting the correct answer and / or if you have any further questions regarding these problems.



Answer by ikleyn(52787)

.
There is entire bunch of lessons covering various types of mixture problems
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Word problems on mixtures for antifreeze solutions
    - Word problems on mixtures for dry substances like coffee beans, nuts, cashew and peanuts
    - Word problems on mixtures for dry substances like candies, dried fruits
    - Word problems on mixtures for dry substances like soil and sand
    - Word problems on mixtures for alloys
    - Typical word problems on mixtures from the archive
    - Advanced mixture problem for three alloys
    - OVERVIEW of lessons on word problems for mixtures
in this site.

Read them and become an expert in solution the mixture word problems.


Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Mixture problems